### v1

# Extreme mass ratio inspirals in dark matter halos: Dynamics and distinguishability of halo models

Abstract
The gravitational wave (GW) signals from extreme mass-ratio inspirals (EMRIs), a key target for the Laser Interferometer Space Antenna (LISA), will be affected in the presence of dark matter (DM) halos. In this paper we explore whether the effects of DM are detectable by LISA within a fully relativistic framework. We model the massive EMRI component as a nonrotating black hole (BH) surrounded by a DM halo. We compute axial and polar GW fluxes for circular orbits at linear order in the mass ratio for DM density profiles with varying mass and compactness. By comparing the phase evolution with vacuum systems, we find that DM halos can induce dephasings of tens to hundreds of radians over a one-year observation period. We demonstrate that even highly diluted DM distributions can significantly affect the emitted waveforms, and that the resulting GW signals can usually be distinguished from each other. While it is important to generalize these findings to more generic orbits and to spinning BHs, our results suggest that LISA could not only reveal the presence of DM halos, but also discriminate between different halo models.
I Introduction
A wide range of astrophysical observations indicate that dark matter (DM) is the most abundant form of matter in the Universe Freese (2009); Navarro et al. (1996); Clowe et al. (2006); Bertone et al. (2005). Understanding the nature of DM is one of the main challenges in modern science, and extensive research efforts are devoted to the exploration of possible interactions between the dark sector and the Standard Model Kahlhoefer (2017); Pérez de los Heros (2020).
In parallel, gravitational wave (GW) observations from coalescing compact objects have opened a novel avenue for probing DM properties and, more generally, the environments in which binaries evolve. The presence of nonvacuum spacetimes can alter both the generation and propagation of GW signals, leaving distinctive imprints on their waveforms. These effects depend on the characteristics of the surrounding environment, which may include baryonic matter in accretion disks, clouds of new fundamental fields, or DM halos Yunes et al. (2011); De Luca et al. (2023); Sberna et al. (2022); Speri et al. (2023); Kavanagh et al. (2020); Macedo et al. (2013); Eda et al. (2013, 2015); Babak et al. (2007); Destounis et al. (2021, 2023); Vicente and Cardoso (2022); Traykova et al. (2021); Sadeghian et al. (2013); Speeney et al. (2022); Bošković et al. (2018); Annulli et al. (2020); Coogan et al. (2022); Cole et al. (2023); Cardoso et al. (2022a, b); Figueiredo et al. (2023); Baumann et al. (2022); Duque et al. (2024a); Barsanti et al. (2023); Duque et al. (2024b); Brito and Shah (2023).
In general, the DM densities required to induce changes in GW signals that may be observable with current or near-future detectors are expected to be several orders of magnitude higher than the average galactic value (in the solar neighborhood, this average density is approximately $\sim 0.01M_{\odot}/{\rm pc}^{3}\sim 0.1{\rm GeV}/{\rm cm}^{3}$). Probing DM through coalescing binaries thus requires a physical mechanism capable of enhancing DM density around isolated compact objects.
A particularly promising scenario involves the formation of DM spikes around black holes (BHs). Both Newtonian and relativistic calculations predict that particle accretion can lead to the development of overdensities, which redistribute to form a cusp. The specific characteristics of these cusps depend on the initial DM profile, with their spatial extent determined by the BH mass Gondolo and Silk (1999); Sadeghian et al. (2013).
The persistence of these overdensities during binary coalescence remains an open question Ullio et al. (2001). Numerical simulations of wave-like DM suggest that overdensities can survive mergers involving comparable-mass binaries, influencing their dynamical evolution Bamber et al. (2023); Zhang et al. (2023); Choudhary et al. (2021); Schive et al. (2014); Aurrekoetxea et al. (2024). In contrast, for models involving heavy DM particles, systems with nearly equal masses are expected to deplete DM spikes, leading to their dissipation during violent or repeated mergers Kavanagh et al. (2018); Bertone et al. (2005); Bertone and Merritt (2005). However, DM overdensities play a crucial role in the dynamics of asymmetric binaries, such as extreme and intermediate mass ratio inspirals (EMRIs/IMRIs).
In particular, EMRIs are promising sources to probe environmental effects through GW observations Cardoso and Maselli (2020). EMRIs consist of a stellar-mass object orbiting a supermassive BH, with mass ratios of $q=m_{p}/M_{\rm BH}\sim\mathcal{O}(10^{-5}-10^{-8})$, where $m_{p}$ and $M_{\rm BH}$ denote the masses of the smaller and larger components, respectively. These systems can spend $\sim 10^{5}$ orbital cycles in the low-frequency regime accessible to LISA Amaro-Seoane et al. (2017), enabling exquisite measurements of source parameters.
Modeling EMRIs in vacuum general relativity (GR) is already a formidable challenge, which is still incomplete after three decades of impressive theoretical developmens Barack and Pound (2019). Extending these calculations to include environmental effects is even more complex, primarily due to two key difficulties: (i) the lack of relativistic solutions describing BHs embedded within matter distributions, and (ii) the intricate couplings between metric and matter variables. As a result, most studies investigating environmental effects on EMRIs rely on approximations, such as post-Newtonian methods or geodesic models that neglect dissipative effects Kocsis et al. (2011); Kavanagh et al. (2020); Coogan et al. (2022); Cole et al. (2023); Tomaselli et al. (2023); Berezhiani et al. (2024).
To address these limitations, a recent research program introduced a relativistic framework based on the Einstein cluster prescription to model the spacetime of BHs surrounded by spherically symmetric density distributions Cardoso et al. (2022a, b). This approach was used to analyze coupled gravitational and fluid perturbations induced by the smaller component of an EMRI when the central BH is embedded in a DM halo. The analysis focused on the Hernquist distribution Hernquist (1990), which allows for closed-form analytic solutions for the background metric. Further advancements extended this formalism by developing a fully numerical framework capable of handling both axial and polar perturbations for general matter distributions surrounding the central BH Figueiredo et al. (2023); Speeney et al. (2024).
In this paper, we extend this research program by investigating the relativistic evolution of EMRIs in a DM halo. We examine how the surrounding matter affects both the adiabatic GW emission and the orbital dynamics for various families of DM density profiles. We quantify deviations from the vacuum case and assess their detectability by LISA using different figures of merit. Our results demonstrate that EMRI observations across a broad mass range can provide strong evidence for the presence of DM spikes near supermassive BHs. Furthermore, we show that the imprints left on GW waveforms by different DM distributions are significant enough to be distinguishable from one another. Unless stated otherwise, throughout the paper we use geometrical units ($G=c=1$).
II Theoretical formalism
In this section we review the formalism developed in Refs. Figueiredo et al. (2023); Speeney et al. (2024) to compute axial and polar perturbations sourced by a massive particle in circular motion around a BH surrounded by a spherically symmetric DM halo.
II.1 Background solution
We consider a static, spherically symmetric spacetime
$ds^{2}=g^{(0)}_{\mu\nu}dx^{\mu}dx^{\nu}=-a(r)dt^{2}+\frac{dr^{2}}{1-\frac{2m(r)}{r}}+r^{2}d\Omega^{2}\ ,$ (1)
where the metric functions $a(r),m(r)$ are solutions of the sourced Einstein field equations
$G_{\mu\nu}^{(0)}=8\pi T^{m(0)}_{\mu\nu}\ .$ (2)
The stress-energy tensor $T^{m(0)}_{\mu\nu}$ is obtained following the Einstein cluster prescription Einstein (1939) and describes the DM content as an anisotropic fluid:
$(T^{m(0)})^{\mu}_{\nu}=\text{diag}(-\rho,0,P_{t},P_{t})\ .$ (3)
The density $\rho$ and the tangential pressure $P_{t}$ depend on the radial coordinate only, and are related through the Bianchi identities:
$P_{t}(r)=\frac{1}{2}\frac{m(r)}{r-2m(r)}\rho(r)\ .$ (4)
From the $tt$ and $rr$ components of Eqs. (2), the mass function $m(r)$ and the metric function $a(r)$ must satisfy
$m^{\prime}(r)=4\pi r^{2}\rho(r)\,,\quad\frac{a^{\prime}(r)}{a(r)}=\frac{2m(r)/r}{r-2m(r)}\ .$ (5)
For a given density profile $\rho(r)$, solving Eqs. (5) fully specifies the background spacetime. We follow the numerical approach developed in Ref. Figueiredo et al. (2023) and consider three families of DM distribution, the Hernquist Hernquist (1990), Navarro-Frenk-White (NFW) Navarro et al. (1997) and Einasto Einasto (1965, 1969) models. The density profile of the first two can be described in terms of a two-parameter semianalytic function
$\rho(r)=\rho_{0}(r/a_{0})^{-\gamma}[1+(r/a_{0})^{\alpha}]^{(\gamma-\beta)/\alpha}\ ,$ (6)
where $a_{0}$ is the scale radius of the halo, and the factor $\rho_{0}$ is proportional to the density, namely $\rho_{0}=2^{(\beta-\gamma)/\alpha}\rho(a_{0})$. The parameters $(\alpha,\beta,\gamma)$ are fixed to $(1,4,1)$ and $(1,3,1)$ for Hernquist and NFW, respectively. Given that the NFW distribution is logarithmically divergent, we introduce a cutoff radius $r_{c}$, such that $M_{\rm H}(r>r_{c})=0$, and assume $r_{c}=5a_{0}$.
The density profile for the Einasto model is given by
$\rho(r)=\rho_{e}{\rm exp}\left\{-d_{n}[(r/r_{e})^{1/n}-1]\right\}\ ,$ (7)
where $n=6$, $d_{n}=53/3$ Graham et al. (2006); Prada et al. (2006), and $\rho_{e}$ is the density at the radius $r_{e}$, that determines a volume containing half of the halo mass. Here we study the evolution of the EMRI by setting $r_{e}=a_{0}$.
Figure 1: Density profiles as a function of distance from a primary BH with mass $M=10^{6}M_{\odot}$. The purple-thick, green-medium, and orange-thin curves correspond to halo compactness values of $\mathcal{C}=10^{-3}$, $\mathcal{C}=10^{-4}$, and $\mathcal{C}=10^{-5}$, respectively. The style of each curve indicates the DM model: Einasto (continuous), Hernquist (dashed), and NFW (dotted). The left and right panels correspond to DM halos with masses $M_{\text{halo}}=10^{2}M_{\text{BH}}$ and $M_{\text{halo}}=10^{3}M_{\text{BH}}$, respectively. Dashed and dotted vertical lines mark the initial orbital radius of a secondary object with mass $m_{p}=10M_{\odot}$ and $m_{p}=50M_{\odot}$, respectively, which evolves for one year around the primary before reaching the innermost stable circular orbit (ISCO). For reference, we also show in black the location of the coordinate radius of the BH horizon and the distance at which the DM profile vanishes.
Finally, for all profiles, we scale the density distribution by a “peak factor” such that $\rho(r)\rightarrow\rho(r)(1-2r_{h}/r)$. This factor is introduced to replicate the overdense cusps with a sharp cutoff at $r=2r_{h}$, with $r_{h}=2M_{\rm BH}$ being the BH horizon, which are expected to result from the adiabatic accretion growth of a nonrotating BH Sadeghian et al. (2013); Gondolo and Silk (1999). We require the density to vanish for $r\leq 2r_{h}$.
Realistic DM profiles, such as those inferred for Milky Way-like galaxies, are expected to have small values of the compactness ${\cal C}\equiv M_{\rm H}/a_{0}\ll 1$. In this regime of compactness, as discussed in Refs. Cardoso et al. (2022a, b), deviations from the Schwarzschild vacuum solution scale linearly with ${\cal C}$. Unlike previous studies, in this work we consider low density, realistic scenarios, focusing on DM halos with ${\cal C}=(10^{-3},10^{-4},10^{-5})$. Moreover, we study binary systems evolving in DM halos with masses $M_{\text{halo}}=(10^{2},10^{3})M_{\text{BH}}$, where $M_{\text{BH}}$ is the mass of the primary BH. Finally, to gauge the dependence of our results on the NFW cutoff radius and on the value of $r_{e}$ for the Einasto model, in Appendix B we provide a complementary analysis for some selected EMRI configurations, assuming different values of $r_{c}$ and $r_{e}$.
In Fig. 1 we show the density profiles of the nine different DM halo configurations analyzed in this work. The scale and shape of the Hernquist and NFW profiles are similar. In contrast, the Einasto model exhibits a steeper fall-off at large distances from the BH, i.e., for $r\gtrsim 10^{5}M_{\odot}$. As a result, the Einasto profile is more compact for a given $M_{\rm halo}$ and $\mathcal{C}$, with densities up to an order of magnitude higher than the Hernquist and NFW distributions at distances $r\sim 10^{3}M_{\rm BH}$.
The dashed and dotted vertical lines in Fig. 1 indicate the initial orbital separation of EMRIs with a secondary mass of $m_{p}=10M_{\odot}$ and $m_{p}=50M_{\odot}$, respectively, evolving over one year until plunge (see Sec. III). This clearly shows that typical EMRI configurations targeted by LISA observations spend the majority of their orbital cycles near the peak of the DM distribution.
II.2 Gravitational and matter perturbations
At the leading dissipative order, the EMRI dynamics is determined by linear perturbations sourced by the light BH orbiting around the massive primary:
$g_{\mu\nu}=g_{\mu\nu}^{(0)}+g_{\mu\nu}^{(1)}\quad\ ,\quad T_{\mu\nu}^{m}=T_{\mu\nu}^{m(0)}+T_{\mu\nu}^{m(1)}\ ,$ (8)
which lead to the perturbed Einstein equations
$G_{\mu\nu}^{(1)}=8\pi T_{\mu\nu}^{m(1)}+8\pi T_{\mu\nu}^{p}\ ,$ (9)
where $T_{\mu\nu}^{p}$ is the stress-energy tensor of the particle (i.e., the secondary compact object). The symmetry of the background allows us to decompose $g^{(1)}_{\mu\nu}$, $T^{m(1)}_{\mu\nu}$ and $T^{p}_{\mu\nu}$ into axial and polar modes, which can be treated independently Regge and Wheeler (1957); Davis et al. (1971). We work in the frequency domain, expanding perturbations in Fourier components, and adopt the Regge-Wheeler-Zerilli gauge.
Matter and tensor modes couple in the polar sector only, with the DM environment affecting axial perturbations only through background quantities. This allows us to obtain a single, second order, differential equation for axial modes that differs from the Schwarzschild case only in the values of $a(r)$ and $m(r)$. The polar sector is more complex and cannot be reduced to a Zerilli-type equation, requiring us to work with a set of coupled ODEs.
In the following we present a brief discussion of the relevant formalism. We refer the reader to Cardoso et al. (2022b); Figueiredo et al. (2023) for more technical details.
II.2.1 Axial perturbations
The axial sector is fully described by two functions, $h_{0}^{\ell m}(r,\omega)$ and $h_{1}^{\ell m}(r,\omega)$, where $\omega$ is the mode frequency, and the indices $(\ell,m)$ specify the multipolar component of the perturbation. Einstein’s equations allow us to express $h_{0}^{\ell m}$ in terms of $h_{1}^{\ell m}$. Moreover, introducing the new variable $\Psi_{\ell m}(r,\omega)=h_{1}^{\ell m}(r,\omega)[a(r)(1-2m(r)/r)]^{1/2}/r$, we obtain a single, second order ODE:
$\frac{d^{2}}{dr_{*}^{2}}\Psi_{\ell m}+[\omega^{2}-V^{\text{ax}}_{\ell}(r)]\Psi_{\ell m}=S_{\ell m}^{\text{ax}}\ ,$ (10)
where the generalized tortoise coordinate $r_{*}$ is defined by
$\frac{dr_{*}}{dr}=a(r)^{-1/2}\left[1-\frac{2m(r)}{r}\right]^{-1/2}\ .$ (11)
The axial potential is
$V_{\ell}^{\text{ax}}(r)=\frac{a(r)}{r^{2}}\left[\ell(\ell+1)-\frac{6m(r)}{r}+m^{\prime}(r)\right]\ ,$ (12)
while the source term $S^{\text{ax}}_{\ell m}$ carries information about the orbit of the secondary.
We solve Eq. (10) using Green’s functions. As a first step we consider the associated homogeneous problem, with the physical solution admitting purely ingoing and outgoing boundary conditions at the horizon ($r=r_{\rm h}$) and infinity $(r\rightarrow\infty)$, respectively. The solution at the boundaries can be expanded as a power series:
$\Psi_{\ell m}^{(\rm in)}=e^{-i\omega r_{*}}\sum_{i=0}^{n_{\text{in}}}\psi^{\rm in}_{i}(r-r_{\text{h}})^{i}\ ,$ (13)
$\Psi^{(\rm out)}_{\ell m}=e^{i\omega r_{*}}\sum_{i=0}^{n_{\text{out}}}\psi^{\rm out}_{i}\frac{1}{r^{i}}\ ,$ (14)
in which we set $n_{\text{out}}=n_{\text{in}}=4$. The values of $(\psi^{\rm in}_{i},\psi^{\rm out}_{i})$ can be found by substituting $(\Psi_{\ell m}^{(\rm in)},\Psi_{\ell m}^{(\rm out)})$ into Eq. (10), expanding in powers of $(r-r_{h})$ and $1/r$ respectively, and solving order-by-order for the coefficients. The leading-order terms can be fixed to $\psi^{\rm in}_{0}=\psi^{\rm out}_{0}=1$.
The full solution at infinity is obtained by integrating the homogeneous solution over the source term:
[…]
where the Wronskian is defined as $W\equiv d\Psi_{\ell m}^{\rm out}/dr_{*}\Psi_{\ell m}^{\rm in}-d\Psi_{\ell m}^{\rm in}/dr_{*}\Psi_{\ell m}^{\rm out}$. The axial energy flux at infinity is then given by
$\dot{E}_{\ell m}^{\infty,\text{ax}}/q^{2}=\frac{1}{8\pi}\frac{(\ell+2)!}{(\ell-2)!}|\Psi^{\rm out}_{\ell m}|^{2}\ .$ (16)
II.2.2 Polar perturbations
Polar modes are described by three metric functions, $K^{\ell m}(r,\omega),H_{0}^{\ell m}(r,\omega)$ and $H_{1}^{\ell m}(r,\omega)$, coupled with the fluid velocity perturbation $W_{\ell m}(r,\omega)$ and the density component $\delta\rho_{\ell m}(r,\omega)$. A barotropic equation of state provides a relation between pressure and density variations via the radial and tangential speed of sound
$c^{2}_{t,r}=\delta p_{t,r}^{\ell m}(r,\omega)/\delta\rho^{\ell m}(r,\omega)\ .$ (17)
Following Cardoso et al. (2022b) we assume the speeds of sound to be constant, although it generally depends on the radial coordinate $r$. This leads to a set of five first-order ODEs, that can be cast in a compact form as
$\ \frac{d\vec{\Phi}_{\ell m}}{dr}-\hat{\alpha}\vec{\Phi}_{\ell m}=\vec{S}_{\ell m}\ ,$ (18)
where $\vec{\Phi}_{\ell m}=\left\{K_{\ell m},H_{0{\ell m}},H_{1{\ell m}},W_{\ell m},\delta\rho_{\ell m}\right\}$, $\vec{S}_{\ell m}=(S_{1\ell m},S_{2\ell m},S_{3\ell m},0,0)$, and the elements of the matrix $\hat{\alpha}$ are shown in the Supplementary Material of Ref. Cardoso et al. (2022b).
The system (18) is solved through a shooting method. We first compute the appropriate boundary conditions of the full inhomogeneous problem, which are purely ingoing at the horizon and purely outgoing at infinity, respectively. At the horizon, we adopt the following ansatz:
$K_{\ell m}^{(\rm in)}=$ $e^{-i\omega r_{*}}\sum_{i}^{n_{\rm in}}k^{\rm in}_{i}(r-r_{\text{h}})^{i}\ ,$ (19)
$H_{0\ell m}^{(\rm in)}=$ $e^{-i\omega r_{*}}\sum_{i}^{n_{\rm in}}h^{\rm in}_{0_{i}}(r-r_{\text{h}})^{i-1}\ ,$ (20)
$H_{1\ell m}^{(\rm in)}=$ $e^{-i\omega r_{*}}\sum_{i}^{n_{\rm in}}h^{\rm in}_{1_{i}}(r-r_{\text{h}})^{i-1}\ ,$ (21)
while, following Ref. Speeney et al. (2024), we fix $W_{\ell m}(r_{\rm h})=\delta\rho_{\ell m}(r_{\rm h})=0$. Similarly, at spatial infinity we introduce a Taylor expansion for the metric functions
$K_{\ell m}^{(\rm out)}=$ $e^{i\omega r_{*}}\sum_{i}^{n_{\rm out}}k^{\rm out}_{i}\frac{1}{r^{i}}\ ,$ (22)
$H_{0{\ell m}}^{(\rm out)}=$ $e^{i\omega r_{*}}\sum_{i}^{n_{\rm out}}h^{\rm out}_{0_{i}}\frac{1}{r^{i-1}}\ ,$ (23)
$H_{1{\ell m}}^{(\rm out)}=$ $e^{i\omega r_{*}}\sum_{i}^{n_{\rm out}}h^{\rm out}_{1_{i}}\frac{1}{r^{i-1}}\ ,$ (24)
and we set $W_{\ell m}(r_{\infty})=\delta\rho_{\ell m}(r_{\infty})=0$. We fix $n_{\rm in}=n_{\rm out}=5$. The coefficients at the horizon and at infinity can be computed with the same iterative procedure discussed for the axial sector. Setting $k^{\rm out}_{0}=1$ completely determines all the coefficients, modulo the shooting parameter $k^{\rm in}_{0}$. For each radius $r$, the solution to Eqs. (18) is found by requiring the perturbations and their derivatives to be continuous at some radius $r_{\infty}$, i.e.,
$\lim_{r\to r_{\infty}}\left[K^{\ell m}_{\text{(sol)}}K^{{}^{\prime}\ell m}_{\text{(out)}}-K^{\prime\ell m}_{\text{(sol)}}K^{\ell m}_{\text{(out)}}\right]=0\ ,$ (25)
where a prime denotes differentiation with respect to $r$, $K_{\text{(sol)}}(r)$ is the numerical solution obtained by direct integration of the ODEs, and $K_{(\rm out)}$ is given by Eq. (22). Finally, the polar energy flux at infinity for each multipolar component is given by
$\dot{E}_{\ell m}^{\infty,\text{pol}}/q^{2}=\lim_{r\to r_{\text{obs}}}\frac{1}{32\pi}\frac{(\ell+2)!}{(\ell-2)!}|K_{\ell m}(r)|^{2}\ .$ (26)
II.3 Geodesics and orbital evolution
Given the background functions $a(r)$ and $m(r)$, we can determine the geodesic properties of particles in the DM halo. The spacetime admits two conserved quantities, corresponding to the energy per unit mass and the specific angular momentum at infinity:
$E_{p}=\left[\frac{r-2m(r)}{r-3m(r)}a(r)\right]^{1/2}_{r=r_{p}}\ ,~{}~{}~{}L_{p}=\left[\frac{m(r)}{r-3m(r)}\right]^{1/2}_{r=r_{p}}\ ,$ (27)
where $r_{p}$ denotes the orbital radius of the particle, and its orbital frequency is given by $\omega_{p}=a(r_{p})L_{p}/r_{p}^{2}E_{p}$. We focus on equatorial orbits, and set $\theta(r=r_{p})=\pi/2$. From $m(r)$ we can also compute the radius of the last stable orbit for massive bodies, $r_{\rm isco}$, which satisfies the following equation:
$r^{2}m^{\prime}(r)+rm(r)-6m(r)^{2}=0\ .$ (28)
Once we have the solution for the background and the GW energy fluxes, we can numerically integrate the equations of motion for the orbital phase and separation:
$\frac{d\Phi}{dt}=\omega_{p}\quad\ ,\quad\frac{dr}{dt}=-\dot{E}\frac{dr}{dE_{\text{orb}}}\ ,$ (29)
where $E_{\text{orb}}$ is the orbital energy of the particle.
We solve Eqs. (29) with initial conditions $\Phi(t=0)=0$ and $r(t=0)=r_{0}$, where $r_{0}$ is chosen such that the EMRI reaches the plunge radius $r_{\rm plunge}$ after time period $T$. We make the conservative choice $r_{\rm plunge}=r_{\rm isco}+0.1M_{\text{BH}}$ because the precision in computing the fluxes decreases for orbits close to $r=r_{\rm isco}$. We will present results for two cases: $T=1$ year and $T=6$ months.
III Numerical setup and detectability
The total energy flux emitted by the binary at each radius is given by
$\dot{E}=\sum_{\ell=2}^{\ell_{\text{max}}}\sum_{m=-\ell}^{\ell}(\dot{E}_{\ell m}^{\infty,\text{ax}}+\dot{E}_{\ell m}^{\infty,\text{pol}})\ .$ (30)
The results presented in the main text have been obtained by setting $\ell_{\text{max}}=2$. In Appendix A, we validate this choice by comparing the results for different values of $\ell_{\text{max}}$. The values of $E_{\ell m}^{\rm ax,pol}$ are computed on a uniform grid of points in the range $r\in[6.0001,14.5001]M_{\text{BH}}$, with step size of $r_{\text{step}}=0.1M_{\text{BH}}$. We solve the equations of motion (29) on this grid for a primary mass $M=10^{6}M_{\odot}$ and two values of the secondary mass: $m_{p}=(10,50)M_{\odot}$. We extract fluxes by setting the numerical infinity in Eqs. (15) and (25) to $r_{\infty}=2\times 10^{7}M_{\rm{BH}}$ when $M_{\rm{halo}}=10^{2}M_{\rm{BH}}$, and $r_{\infty}=2\times 10^{8}M_{\rm{BH}}$ when $M_{\rm{halo}}=10^{3}M_{\rm{BH}}$. These values correspond to $2a_{0_{\rm{max}}}$, where $a_{0_{\rm{max}}}$ is the maximum value of $a_{0}$ among the three different choices used to define the compactness for a fixed value of $M_{\rm{halo}}$ (we have also checked that changes in $r_{\infty}$ by one order of magnitude do not significantly affect the GW fluxes). Finally, to ensure a proper comparison with vacuum fluxes (and consequently with GW phases and waveforms), we compute the vacuum fluxes at the same numerical infinity.
Figure 2: Top row: Ratio of GW fluxes emitted in vacuum to those in the presence of matter, plotted as a function of the secondary’s orbital radius. The primary BH mass is $M=10^{6}M_{\odot}$, the DM halo mass is $M_{\text{halo}}=10^{2}M_{\text{BH}}$, and we consider two possible values of the secondary mass: $m_{P}=(10,50)M_{\odot}$. The purple solid, green dashed, and orange dotted lines represent halos with compactness values of $\mathcal{C}=(10^{-3},10^{-4},10^{-5})$. Markers along each curve indicate the initial orbital radius of the secondary, which evolves over either one year (filled markers) or six months (hollow markers) until plunge (see legend). Bottom row: Quadrupolar dephasing accumulated over one year (filled markers) and six months (hollow markers) during the evolution of the EMRIs. Labels on the top axes indicate the accumulated number of cycles in the vacuum case. The three different panels, from left to right, correspond to different DM density profiles.
Figure 3: Same as Fig. 2, but assuming a DM halo with $\text{M}_{\text{halo}}=10^{3}\text{M}_{\text{BH}}$.
To assess the environment detectability we consider two figures of merit. We first study the evolution in phase of the GW signal, computing the phase difference between binaries in vacuum and in a DM background, $\Delta\Phi=\Phi_{v}-\Phi_{m}$, after a given observation time $T$. A value of $\Delta\Phi$ larger than a certain threshold provides an indication that the environment can be detected by LISA. Following Bonga et al. (2019), we assume the threshold to be $\Delta\bar{\Phi}=0.1$ radians, which roughly corresponds to a resolvable phase for a system with signal-to-noise ratio of $\sim 30$.
A more quantitative and accurate assessment can be made by evaluating waveform changes weighted by the detector sensitivity. To this aim we compute the faithfulness ${\cal F}$ between GW signals from EMRIs in vacuum, and those surrounded by the DM distribution:
$\mathcal{F}\left[h_{m},h_{v}\right]=\max_{\{t_{c},\phi_{c}\}}\frac{{\langle h_{m}|h_{v}\rangle}}{\sqrt{{\langle h_{m}|h_{m}\rangle}{\langle h_{v}|h_{v}\rangle}}}\ ,$ (31)
where
[…]
Here $S_{n}(f)$ is the LISA noise power spectral density, including the confusion noise produced by unresolved galactic white dwarf binaries Robson et al. (2019). The frequency domain waveforms $\tilde{h}_{m,v}(f)$ are computed by applying a discrete Fourier transform to the time domain plus and cross signal polarizations. These are derived using the quadrupolar formula Huerta and Gair (2011); Barack and Cutler (2004), i.e.,
$h_{+}=(\ddot{I}_{11}-\ddot{I}_{22})/2\ ,\quad h_{\times}=\ddot{I}_{12}\ ,$ (33)
where a dot represents a derivative with respect to the coordinate time, the quadrupole moment tensor is defined as $I_{ij}=2m_{p}/d_{\rm L}z^{i}(t)z^{j}(t)$, $z(t)$ is the geodesic of the secondary expressed in Cartesian spatial coordinates, and $d_{\rm L}$ is the source luminosity distance. We refer the reader to the method section of Maselli et al. (2022) for further technical details on the waveform generation, and on the numerical implementation of Eq. (31). Finally, we fix $f_{\text{min}}=10^{-4}$ Hz in the integral of Eq. (32), while $f_{\text{max}}$ corresponds to the Nyquist frequency of the signal. We assume that two waveform models are distinguishable if the faithfulness is smaller than the critical value ¯$\bar{{\cal F}}\simeq 0.92$, corresponding to the estimated distinguishability threshold for a template depending on ten parameters, observed with a signal-to-noise ratio of $\sim 10$ Chatziioannou et al. (2017).
IV Results
We assess the detectability of three families of DM halos: Hernquist, NFW, and Einasto models. We assume halo compactness values of $\mathcal{C}=(10^{-3},10^{-4},10^{-5})$, which are lower than those in previous studies and provide a more realistic description of DM densities within overdensities extending up to $10^{7}$ gravitational radii from the central BH.
In Fig. 2 (top row) we show the ratio of matter to vacuum fluxes for EMRIs evolving in DM halos with $M_{\rm halo}=10^{2}M_{\odot}$ as a function of the secondary’s orbital radius. Filled (hollow) markers represent systems evolving for one year (six months) in the LISA band before plunge, respectively, with each marker positioned at the initial inspiral radius. Deviations from the Schwarzschild case can reach up to $10^{-3}$, and generally increase (if only mildly) as a function of the initial orbital separation.
In many cases, fluxes in matter are lower than those in vacuum for orbital separations close to the ISCO. As shown in Refs. Cardoso et al. (2022b), this reduction can be partially explained in terms of the gravitational redshift induced at the leading order in the halo compactness by the extended distribution of DM. However, $\dot{E}_{v}/\dot{E}_{m}$ has a local minimum at $r\sim 7M_{\rm BH}$ for secondaries closer to the BH. We attribute this to the DM radial distribution, which peaks just below $r\sim 10M_{\rm BH}$ (see Fig. 1). In regions of higher DM density, nonlinear contributions from $M_{\rm halo}$ and $a_{0}$ may become more relevant, counteracting redshift effects. The ratio $\dot{E}_{v}/\dot{E}_{m}$ tend to grow again as $r$ approaches $r_{\rm ISCO}$, where the density $\rho(r)$ decreases.
While the redshift scaling is the primary effect for axial fluxes, the fluxes in the polar sector can also be affected by couplings between the metric and fluid modes. However, we have verified that for the values of ${\cal C}$ considered here such couplings have a very small effect. We have numerically integrated the ODE system (18) by artificially setting matter perturbations to zero, and we find values of $\dot{E}_{m}$ almost indistinguishable from those in Fig. 2. For small values of the halo compactness, changes in the EMRI evolution are primarily driven by the non-Schwarzschild background in which the secondary moves.
The flux differences accumulate throughout the inspiral due to the large number of orbits and they affect the binary’s phase evolution, as shown in the bottom panels of Fig. 2. For all the configurations analyzed here the dephasing lies well above the detection threshold, reaching values as high as $\Delta\Phi\sim 10^{3}$ radians over one year. As expected, the dephasing increases with ${\cal C}$, i.e., for more compact DM distributions. We find that a linear fit $\Delta\Phi=a+b{\cal C}$ describe all our data with an accuracy better than $1\%$, regardless of the halo model and binary configuration. EMRIs with lighter secondaries accumulate a larger number of cycles (as shown on the top x-axis of each panel), and therefore they have the largest dephasings $\Delta\Phi$.
Comparing different halo models, Hernquist and NFW yield similar results, while the Einasto distribution leads to the largest values of $\Delta\Phi$. This is because the axial and polar fluxes are affected by the local value of $\rho(r)$ appearing in Eqs. (10) and (18). The radial profiles in Fig. 1 show that the densities of Hernquist and the NFW models are very close – within a factor two in the region in which the secondary inspirals. In contrast, the Einasto distribution has higher densities, exceeding Hernquist and NFW by a factor $\sim 10-30$ within a distance of $\sim 10M_{\rm BH}$ from the ISCO.
Figure 4: Faithfulness density matrices for different DM profiles (Hernquist, NFW, and Einasto) of mass $M_{\text{halo}}=10^{2}M_{\text{BH}}$, compared to the vacuum case and among each other. The three panels correspond to halo compactness values of $\mathcal{C}=(10^{-3},10^{-4},10^{-5})$. The upper and lower diagonals of each matrix display faithfulness values computed for systems with $(M_{\text{BH}},m_{p})=(10^{6},10)M_{\odot}$ and $(M_{\text{BH}},m_{p})=(10^{6},50)M_{\odot}$, respectively. Large (small) numbers in each cell indicate faithfulness values computed for a one-year (six-month) evolution, respectively.
Figure 5: Same as Fig. 4, but for a halo mass of $M_{\text{halo}}=10^{3}M_{\text{BH}}$.
In Fig. 3 we repeat the dephasing analysis for halos of $M_{\rm halo}=10^{3}M_{\odot}$. The results are qualitatively and quantitatively similar to those for lighter DM distributions. This confirms that the phase evolution in the halo is primarily driven by its compactness, rather than the total halo mass and size.
To further support this conclusion, we have computed the GW fluxes and evolved EMRIs assuming halos with the same compactness as the models discussed so far. However, we considered different combinations of $M$ and $a_{0}$, specifically by setting $M_{H}=M_{\rm BH}$ and $a_{0}=(10^{3},10^{4},10^{5})M_{H}$. When comparing the dephasing relative to vacuum evolutions, we found minor differences (of the order of a few percent) only for the largest valus of the compactness, ${\cal C}=10^{-3}$. Halos with lower compactness did not lead to any significant changes compared to the EMRIs previously analyzed.
While promising, these results do not account for parameter correlations in the waveform model or the impact of LISA noise. To assess the relevance of halo effects in realistic GW observations, we have computed the faithfulness ${\cal F}$ defined in Eq. (31) between waveforms in matter and vacuum for the EMRI configurations analyzed above. Our results are summarized in the matrix diagrams of Figs. 4 and 5, where we list ${\cal F}$ for six months and one year of observation, $M_{\rm halo}=(10^{2},10^{3})M_{\odot}$, and for the three compactness values considered above. For both $m_{p}=10M_{\odot}$ and $m_{p}=50M_{\odot}$, GW signals from binaries evolving in DM environments differ significantly from vacuum, with ${\cal F}\ll\bar{{\cal F}}$ (where ¯$\bar{{\cal F}}=0.92$ is the distinguishability threshold discussed in Sec. III) even in the case of diluted distributions with ${\cal C}=10^{-5}$.
The off-diagonal blocks in the matrices of Figs. 4 and 5 show the faithfulness computed by comparing different halo profiles against each other. They suggest that LISA observations may not only distinguish matter effects from vacuum, but also discriminate between different radial density profiles. Even in the case of signals with similar halo models, like Hernquist and NFW, we obtain ${\cal F}<1$ in most cases, with the exception of halos with ${\cal C}=10^{-5}$.
The faithfulness is typically mildly affected as we vary the secondary mass from $m_{P}=10M_{\odot}$ to $m_{P}=50M_{\odot}$. In general we would expect the faithfulness for one-year observations to be smaller than the for six-month observations, but this is not always the case. This may be due to subtle variations in the waveform phase between the different matter distributions, which can sometimes lead to larger overlaps when the observation time is longer.
By comparing Figs. 4 and 5, we see that these broad conclusions are insensitive to variations in the halo mass.
As a final remark, we want to emphasize the importance of using a fully relativistic description to model EMRI dynamics within DM halos. To demonstrate this, we have computed the phase difference between binaries in which the orbital elements in Eqs. (29) evolve with two different schemes: one based on the $\ell=m=2$ component of the GW flux in Eq. (30), and the other using the quadrupolar post-Newtonian (PN) prescription from Ref. Speeney et al. (2022):
$\dot{E}_{\rm GW}^{\infty}=\frac{32}{5}(M_{\rm BH}\omega_{p})^{10/3}\left[1+\frac{4}{3}\delta(r)\right]\ ,$ (34)
where $\delta(r)=M_{\rm halo}(r)/M_{\rm BH}$, with $M_{\rm halo}(r)$ representing the mass of the halo enclosed within a sphere of radius $r$. In Fig. 6 we show the accumulated dephasing as a function of time until plunge, for a one-year observation of EMRIs with a primary mass of $M=10^{6}M_{\odot}$ and a secondary of $m_{p}=(10,50)M_{\odot}$. We consider a DM halo modeled by the Einasto profile with two different values of $\mathcal{C}$, but we have checked that other DM profiles yield similar results. Even when $\mathcal{C}=10^{-5}$ the PN approximation quickly becomes inadequate, with the accumulated phase difference $\Delta\Phi$ reaching hundreds of radians within just a few months of evolution.
Figure 6: Quadrupolar dephasing accumulated over one year of evolution for EMRIs with a primary mass of $M=10^{6}M_{\odot}$ and a secondary of $m_{p}=10M_{\odot}$ (filled markers) or $m_{p}=50M_{\odot}$ (hollow markers). The binaries evolve within a dark matter halo described by the Einasto profile, with a halo mass of $M_{\text{halo}}=10^{2}M_{\text{BH}}$ and compactness values of $\mathcal{C}=10^{-3}$ (purple round markers) and $\mathcal{C}=10^{-5}$ (orange square markers). The dephasing is computed by comparing EMRIs evolving under two different prescriptions for the GW energy flux. In one case, we use the $l=2,m=2$ component of the relativistic flux from Eq. (26). In the other, we apply the quadrupolar PN expression given by Eq. (34).
V Conclusions
In this work, we have studied the relativistic dynamics of EMRIs evolving within DM distributions. We computed the gravitational and fluid perturbations induced by a light compact object orbiting a nonrotating BH embedded in a spherically symmetric halo, evaluating the GW energy flux and the binary phase evolution for different halo models. We have analyzed three widely used DM models (the Hernquist, Navarro-Frenk-White, and Einasto distributions) with different total mass and characteristic scale and different binary masses, assessing the ability of LISA observations to detect deviations from vacuum predictions.
We have found that EMRIs evolving in DM environments can accumulate phase differences of hundreds of radians compared to their vacuum counterparts over a one-year evolution in the LISA band. For a given DM profile, this dephasing is primarily determined by the halo compactness, scaling linearly with ${\cal C}$. Lighter secondaries experience larger phase deviations due to their slower inspiral and higher number of accumulated cycles. For a fixed compactness, the phase difference $\Delta\Phi$ depends on the maximum local density in the orbital region where the secondary evolves. Even in diluted configurations with ${\cal C}=10^{-5}$, EMRIs with a secondary mass of either $10M_{\odot}$ or $50M_{\odot}$ would experience a phase shift of approximately 10 radians after roughly two months of observation.
These findings are further supported by a refined analysis we performed to assess the distinguishability of GW signals weighted by the LISA sensitivity curve, in which we computed the faithfulness between waveforms emitted in vacuum and in the presence of DM. We find that six months of observation are already sufficient to confidently discriminate between EMRIs in vacuum and EMRIs evolving in DM profiles. More interestingly, we find that LISA can also differentiate among DM profiles with different radial density distributions.
A complete assessment of the ability of future observations to constrain DM models, however, requires dedicated parameter estimation techniques. The waveform models computed in this work can be readily incorporated into LISA data analysis pipelines such as Fast EMRI Waveforms Katz et al. (2021). However, performing full Bayesian (or even simple Fisher matrix) parameter estimation calculations for EMRIs remains particularly challenging because of the extreme accuracy required in GW flux calculations. We plan to address this in a follow-up study. The BH solutions we derived can be extended to density-pressure profiles not necessarily associated with DM halos, such as accretion disks. Finally, recent studies have emphasized the importance of the BH spin in environmental effects on GW signals Dyson et al. (2025). As a natural extension of this work, we will study EMRIs evolving in a Kerr BH background.
Figure 7: Faithfulness density matrices for different DM profiles (Hernquist, NFW, and Einasto) compared to the vacuum case and among each other. The three panels correspond to halo compactness values of $\mathcal{C}=(10^{-3},10^{-4},10^{-5})$, assuming a halo mass of $M_{\text{halo}}=10^{2}M_{\text{BH}}$. The upper and lower diagonals of each matrix display faithfulness values computed for systems with $(M_{\text{BH}},m_{p})=(10^{6},10)M_{\odot}$ and $(M_{\text{BH}},m_{p})=(10^{6},50)M_{\odot}$, respectively. Large numbers in each cell indicate faithfulness values computed for a one-year evolution, while smaller numbers correspond to results obtained for a six-month evolution. At variance with the results shown in Fig. 4, in the first and second row we assume $(r_{c}=a_{0},r_{e}=a_{0})$ and $(r_{c}=5a_{0},r_{e}=a_{0}/2)$ for the NFW and the Einasto profiles, respectively.
Acknowledgements.
N.S. and E.B. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation. A.M. and S.G. acknowledge financial support from MUR PRIN Grants No. 2022-Z9X4XS and No. 2020KB33TP. The Authors acknowledge support from the ITA-USA Science and Technology Cooperation program, supported by the Ministry of Foreign Affairs of Italy (MAECI), grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (https://www.arch.jhu.edu/), which is supported by the NSF Grant No. OAC-1920103.
[…]
profile $[M_{\odot}]$ [year] $\ell_{\rm max}=2$ $\ell_{\rm max}=3$ $\ell_{\rm max}=4$ $\ell_{\rm max}=2$ $\ell_{\rm max}=3$ $\ell_{\rm max}=4$
Hernquist 10 1 436 416 412 0.41 0.41 0.41
0.5 180 170 169 0.42 0.42 0.41
50 1 269 256 255 0.39 0.40 0.40
0.5 100 96 96 0.42 0.41 0.41
NFW 10 1 386 367 364 0.41 0.41 0.42
0.5 159 151 150 0.41 0.41 0.42
50 1 237 226 225 0.40 0.40 0.40
0.5 89 85 84 0.41 0.42 0.41
Einasto 10 1 1360 1310 1303 0.41 0.41 0.41
0.5 557 535 532 0.41 0.41 0.41
50 1 846 816 812 0.39 0.39 0.39
0.5 320 309 308 0.40 0.40 0.40
Table 1: GW dephasing $\Delta\Phi$ and faithfulness ${\cal F}$ with respect to vacuum computed for EMRIs with a primary of $M=10^{6}M_{\odot}$ and a different secondary masses ($m_{p}=10M_{\odot}$ or $m_{p}=50M_{\odot}$), evolving for either one year or six months in different DM profiles. We assume halo masses of $\rm{M}_{\rm{halo}}=10^{2}\rm{M}_{\odot}$ and a compactness of ${\cal C}=10^{-3}$. The dephasing and faithfulness are computed using the GW fluxes in Eq. (30), truncating the sum over $\ell$ at $\ell=\ell_{\rm max}$. The values for $\ell_{\rm max}=2$ correspond to the results discussed in Sec. IV.
Appendix A Dependence on $\ell_{\text{max}}$
In the main text we computed GW fluxes fixing $\ell_{\rm max}$ in Eq. (30) to the leading ($\ell=2$) contribution. In this appendix we assess the impact of higher-order multipoles in the calculations of the dephasing and faithfulness. We compare the results presented in Sec. IV with the values of $\Delta\Phi$ and ${\cal F}$ obtained by including higher multipoles ($\ell_{\mathrm{max}}=2,3,4$) across all three density profiles and for different values of the secondary mass. We focus on EMRIs with halo mass $\mathrm{M}_{\mathrm{halo}}=10^{2}M_{\mathrm{BH}}$ and compactness $\mathcal{C}=10^{-3}$, as these parameters yield the largest deviations from the vacuum scenario. The results are presented in Table 1. While increasing $\ell_{\rm max}$ reduces the dephasing by $\sim 5\%$ relative to the baseline configuration, this effect tends to saturate for all halo models and EMRI configurations. More importantly, the distinguishability between signals in vacuum and matter environments remains unaffected. The faithfulness hardly changes at all when we consider different values of $\ell_{\rm max}$. This confirms the robustness of our main findings.
Appendix B The effect of halo cutoff parameters on the DM detectability
To assess the impact of the NFW cutoff radius and of the choice of $r_{e}$ in Eq. (2) for the Einasto model, we compute the faithfulness with respect to vacuum signals for the same EMRI configurations shown in Fig. 4, varying both $r_{c}$ and $r_{e}$. The first and second rows of Fig. 7 display the values of ${\cal F}$ for a DM halo of $M=100M_{\odot}$, assuming $(r_{c}=a_{0},r_{e}=a_{0})$ and $(r_{c}=5a_{0},r_{e}=a_{0}/2)$, respectively. The values of ${\cal F}$, whether computed against vacuum templates or between different DM models, remain consistent with those in Fig. 4, where we assumed $(r_{c}=5a_{0},r_{e}=a_{0})$. This confirms that the distinguishability of halo signals is largely insensitive to the mass and scale of the distribution, while it primarily depends on the overall halo compactness.